Project

General

Profile

Actions

Issues

Filters

Apply Clear

# Tracker Status Priority Subject Assignee Updated
936 Task In Progress Normal GEM polarimeter cabling Holly Szumila-Vance 02/21/2024 10:11 AM Actions
934 Task New Normal BB Shower VTP firmware 02/05/2024 11:14 AM Actions
933 Task In Progress High HCAL trigger firmware Ben Raydo 03/25/2024 11:20 AM Actions
932 Task In Progress Normal RF time check Alexandre Camsonne 02/28/2024 04:30 PM Actions
666 Task New Normal Cosmics 06/01/2022 06:36 AM Actions
665 Task New Normal VTP readout 06/01/2022 06:36 AM Actions
664 Task New Normal Connect fibers Alexandre Camsonne 06/01/2022 06:35 AM Actions
663 Task New Normal Timing diagram Alexandre Camsonne 06/01/2022 06:33 AM Actions
662 Task New Normal Add SBS VTP 06/01/2022 06:31 AM Actions
661 Task New Normal Fibers for MPD patch 06/01/2022 06:31 AM Actions
660 Task New Normal VETROC readout 06/01/2022 06:14 AM Actions
659 Task New Normal VETROC install 06/01/2022 06:14 AM Actions
658 Task New Normal SD recabling 06/01/2022 06:10 AM Actions
657 Task New Normal Test on GMn data 06/01/2022 06:05 AM Actions
656 Task New Normal Training 06/01/2022 06:05 AM Actions
655 Task New Normal Setup 06/01/2022 06:05 AM Actions
654 Task New Normal Stream Merger 06/01/2022 05:59 AM Actions
653 Task New Normal EPICS scripts for 3 streams readout Alexandre Camsonne 06/01/2022 05:59 AM Actions
651 Task New Normal BPM/Raster weldment 07/15/2021 08:50 PM Actions
650 Task New Normal Install 1 more VXS crate weldment 07/15/2021 08:50 PM Actions
649 Task New Normal Router installation 07/15/2021 08:44 PM Actions
619 Bug New Normal Scalers rate for triggers 05/27/2021 01:42 PM Actions
618 Feature New Normal Re-configure F1 clock to use TI clock 03/29/2021 04:16 PM Actions
617 Feature New Normal Re-configure F1 clock to use TI clock 03/29/2021 07:53 PM Actions
598 Bug New Normal Read out error with crate1 MPD12 10/23/2020 10:24 AM Actions
(51-75/90) Per page: 25, 50, 100

Also available in: Atom CSV PDF